ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 35, 81 Raikov, George, Bulgarian Academy of Sciences, Sofia, Bulgaria Eigenvalue asymptotics for the Dirac operator in strong constant magnetic fields The Dirac operator with constant magnetic field and electric potential decaying at infinity is considered. The asymptotic behaviour of the discrete spectrum for large magnetic-field coupling constants is investigated.